OpenStudy (anonymous):

please helpp if f(x) is 1-1 and if f(x) is never zero can any be said about h(x) = 1/f(x)? is it also 1-1? Give reasons for your answer?

OpenStudy (precal):

is that the actual question? Can you state the question given in your class? Seems like it is miswritten or something. What level of math is this? just curious so I know how to answer it, if I can

OpenStudy (anonymous):

its first year calculus and that is what is written in the book

OpenStudy (kinggeorge):

I do believe that would also be 1-1. If it were 0, we would run into a problem dividing by 0, but it isn't, and since \(f(x)\) takes unique values for unique \(x\)'s, \(1/f(x)\) must also take unique values.

OpenStudy (anonymous):

hmm so just again 1-1 basically means it passes the horizantal and the vertical line test?

OpenStudy (kinggeorge):

1-1 means it passes the horizontal line test. However, it must also pass the vertical line test since that's required to be a function.

OpenStudy (anonymous):

could we also prove this by defining a 1-1 function saying let us pick two arbitrary points say a,b and h(a) = 1/f(a) and h(b) = 1/f(b) therefore 1/f(a)=1/f(b) and that implies f(b)=f(a) and that implies b=a and that implies h is a 1-1 function ??

OpenStudy (kinggeorge):

You would prove it by saying that: Suppose h(a)=1/f(a)=h(b)=1/f(b). Then 1/f(a)=1/f(b) so f(a)=f(b). Since f is 1-1, this implies that a=b. Hence, h is also 1-1.

OpenStudy (anonymous):

perfect thanks a lot

OpenStudy (anonymous):

could you also help me in this other question please i am going to post please